OpenStudy (anonymous):

Stuck on this one multiply and simplify √125*√250 =√25*5*25*10 =5*5√5*10 =25√50 Doesnt sound right

OpenStudy (king):

its rite but 25sqrt50 can be written as 25sqrt 5*5*2 that is 125sqrt2 \[125\sqrt{2}\]

OpenStudy (anonymous):

ah i see. 25*25√2 right

OpenStudy (anonymous):

I was just missing one more step than

OpenStudy (king):

no!!it is not 25*25

OpenStudy (anonymous):

i mean 25*5

OpenStudy (anonymous):

cause of the square of 25

OpenStudy (king):

\[25\sqrt{5\times5\times2}\] \[25\times5\sqrt{2}\]

OpenStudy (king):

yeah \[125\sqrt{?}\]

OpenStudy (king):

sry\[125\sqrt{2}\]

OpenStudy (anonymous):

ok makes sense. I was just looking at it wrong

OpenStudy (phi):

with these problems, factor everything into prime factors and collect terms. Then look for primes with an even power to pull out of the square root.